Centre de Données astronomiques de Strasbourg

———

Reference Codes in SIMBAD

25-May-1994 May 25, 1994

Equipe Mise à Jour du CDS

Codes Références dans Simbad  (au 25-May-1994)  1

A&A... Astron. Astrophys. AAA... Astronomy and Astrophysics Abstracts,

Heidelberg

AAfz.. Astrometria Astrofizika AAHam. Astron. Abh. Hamburg. Sternw. AAONw. Anglo-Australian Observatory Epping -

Newsletter

AAOPr. Anglo-Australian Obs. Prepr. A&ARv. Astron. Astrophys. Rev. A&AS.. Astron. Astrophys., Suppl. Ser. AASPP. Astron. Astrophys. Serie, Ed. Pachart

Publishing House Tucson

A&ASS. Astron. Astrophys., Special Suppl. Ser. AbaOB. Abastumanskaya Astrofiz. Obs., Byull. AbhKP. Abh. Konigl. Preuss. Akad. Wis-

senschaften Jahre 1906,92, Berlin,1907 (XXAF)

AcA...   Acta Astron.

AcApS. Acta Astrophys. Sinica AcAS.. ACta Astronomica Supplementa AcASn. Acta Astron. Sinica AcC... Acta Cosmologica

ACiCh.   Astron. Cir.

ACMan. Astronomical Contributions from the

University of Manchester

AcMPh. Acta Univ. Carol. Mat. Phys. AcPhA. Acta Phys. Austriaca ADS... Carnegie Inst. Washington D.C. Publ.

417

ADUrb. University of Illinois. Astronomy De-

partment, Urbana Illinois

AExpr. Astron. Express AFGL.. AFGL-TR-0208 Environemental

Research papers,576,1976 (XXGI)

AFOEV. Bull. Assoc. Fr. Obs. Etoiles Variables

Afz...   Astrofizika

AGAb.. Astron. Gesellschaft Abstract Ser. AGK2.. Hamburger Sternwarte

Bergedorf,Bonn,1958 (XXCG)

AISAO. Astrofizik. Issledovanija, special Obs. AISof. Astrofisik. Issledovanija - Sofia

AJ....   Astron. J.

AJa... Astronomischer Jahresbericht AL.... Astron. Lett. AN.... Astron. Nachr. AnAp.. Ann. Astrophys. AnBes. Annales de l’Observatoire de Besancon AnBos. Annals of the Bosscha Observatory Lem-

bang (Java) Indonesia

AnCap. Annals Cape Obs. AnDea. Ann. Dearborn Obs. ANErg. Ergaenzungshefte zu den Astronomis-

chen Nachrichten AnHar. Ann. Astron. Obs. Harvard Coll. AnLei. Ann. Sternw. Leiden ANote. Astronomical Notes - Univer-

sity of Gothenburg, Sweden, Section of Astronomy

AnPar. Ann. Obs. Paris AnTok. Ann. Tokyo Astron. Obs. AnTou. Ann. de l’Observ. Astr. et Meteo. de

Toulouse

AnWie. Ann. Univ. Sternw. Wien ANZi.. Beobachtungs-Zirkular der Astronomis-

chen Nachrichten

Ap....   Astrophys.

ApJ... Astrophys. J. ApJS.. Astrophys. J., Suppl. Ser. ApL... Astrophys. Lett. ApNr.. Astrophys. Norvegica Ap&SS. Astrophys. Space Sci. A&R... Astron. Raumfahrt

AR....   Astron. Rep.

ArA... Arkiv for Astron. ARA&A. Annual Rev. Astron. Astrophys. ARKod. Annual Report of the Kodaikanal

Observatory

ArMAF. Ark. Mat. Astr. Fys. ArmOL. Armagh Observatory - Leaflet

ArS...   Arch. Sci.

ASSU.. Annual Scientific Supplement to Urania

Aster.   Aster

Astro.   Astronomy

ATi... Astron. Tidsskr. AtlSV. Specola Astronomica Vaticana - Atlas

Stellarum Variabilium

AtlVS. Atlas Poiskovykh Kart Peremennykh

Zvezd

ATsir. Astron. Tsirk. AtSV.. Charts for Southern Variables AuJPA. Australian J. Phys. Astrophys. Suppl. AuJPh. Australian J. Phys. AuJPS. Australian J. Phys. Suppl. AVSOA. American Association of Variable Stars

Observers - Abstracts

AVSOM. American Ass. Var. Stars Obs.

monograph

AVSOQ. American Association of Variable Stars

Observers - Quarterly Reports

AVSOR. Am. Assoc. Variable Stars Obs. Rep AVSVB. Associazione

Veneta Osservatori di Stelle Variabili - Bulletin

AZh...   Astron. Zu.

BAAA.. Bol. de la Asociacion Argent. de Astron. BAAS.. Bull. American Astron. Soc. BAICz. Bull. Astron. Inst. Czech. BaltA. Baltic Astronomy BamKV. Bamberg Kleine Veroeffentlichungen der

Remeis-Sternwarte BamVe. Bamberg Veroeffentlichungen der

Remeis-Sternwarte

BAN... Bull. Astron. Inst. Netherlands BANS.. Bulletin of the Astronomical Institutes

of the Netherlands, Supplement Series

BAORB. Bull. Astron. Obs. R. Belgique

Bar1K.   Barcelone

BARB.. Bull. Cl. Sci., Acad. R. Belg. BASI.. Bull. Astron. Soc. India BAVSM. Berliner Arbeitsgemeinschaft fuer Ver-

aenderliche Sterne - Mitteilungen

BAVSO. Bulletin of the American Association of

Variable Stars Observers

BAVSR. BAV Rundbrief

BCFHT.   Bull. CFH

BCrAO. Bull. Crimean Astrophys. Obs. BDCat. Eds Marcus A. and E. Weber’s Verlag

Bonn

BDSud. Eds Marcus and Weber’s Verlag, Bonn BDus.. Bjull. Inst. Astrofizikii BeiRe. Reprints, Beijing Astronomical Observa-

tory, Academia Sinica

Bes1K. Observatoire de Besancon BHarO. Harvard Bull. BICDS. Bull. Inf. Centre Donnees Stellaires

BID1..   bidon1

BIMAF.   ?????????

BITA.. Bjulleten Inst. Teoret. Astron. BITon. Bol. Inst. Tonantzintla BKobO. Bulletin of the Kobe Marine Observa-

tory, Kobe, Japan

BMai.. Bulletin de la Station

Astrophotographique de Mainterne

BOBeo. Bull. Obs. Astron. Beograd BOPul. Bull. Obs. Poulkovo BOTor. Biul. Obs. Astron. Uniw. N. Kopernika

Toruniu

BOTT.. Bol. Obs. Tonantz. Tacub. BPM... The Observatory, Univ. Minnesota,

Minneapolis,1963 (XXDT)

BS.... Bulletin Signaletique - Centre National

de la Recherche Scientifique

BSD... Hamburger Sternwarte Bergedorf BSRSL. Bull. Soc. R. Sci. Liege BSVIt. Bolletino della Societa dei Variabilisti

Italiani

BYam.. Bulletin of the Yamagata University, Ya-

magata, Japan

ByuPr. Byurakanskaya Astrofizicheskaya Obser-

vatoriya - Preprint

ByuRe. Byurakan Astrophys. Obs. CAFOE. Circulaire de l’Association Francaise

d’Observateurs d’Etoiles Variables

CamRe. Cambridge Observatories - Reprints CarOB. Carter Observatory, Wellington, New

Zealand - Astronomical Bulletins CarRe. Carter Obs., Repr. CDT... Connaissance des Temps C&E... Ciel et Espace CfAPr. Cent. Astrophys., Prepr. Ser. CFC... London (on magnetic tape),1939

(XXAM)

CfMil. Conferenze dell’Osservatorio

Astronomico di Milano-Merate

CGCG.. California Inst. Techn. ChA&A. Chinese Astron. CiBAA. Circular of the British Astronomical

Association

CiUO.. Union Obs. Circ. CoAnk. Commun. Astron. Dep., Univ. Ankara CoAp.. Comments on Astrophys. CoArm. Contributions from the Armagh

Observatory

CoAsi. Contr. Oss. Astrof. Padova in Asiago CoBos. Contr. from the Bosscha Observ.

Lembang.

CoBrn. Contributions of the Public Observatory

and Planetarium in Brno

CoCam. Contr. Cambridge Obs. CoCoi. Comun. Obs. Astron. Univ. Coimbra CoDAO. Contr. Dominion Astrophys. Obs.

Victoria

CoDDO. Com. from the David Dunlap Observ. CoDun. Contrib. Dunsink Obs. Coel.. Coelum - Periodico Bimestrale per la Di-

vulgazione dell’Astronomia

CoFra. Contr. Cons. Nazion. Ric. Frascati CoIAB. Contributions de l’Institut

d’Astrophysique de Paris. Serie B.

CoIAP. Contr. de l’Inst. d’Astr. de Paris CoKit. Contributions from the Kitt Peak Na-

tional Observatory

CoKon. Comm. Konkoly Obs. CoKwa. Contr. Astroph. Kwasan Obs. Univ.

Kyoto

CoLic. Contr. Lick Obs. CoLon. Comm. Univ. London Obs. CoLPL. Comm. Lunar Plan. Lab. CoMil. Contr. dell’Osserv. Astron. di Milano-

Merate

CoMtW. Contributions from the Mount Wilson

Observatory

CoORB. Com. de l’Observ. Royal de Belgique CoOxf. Communications from the University

Observatory, Oxford

Cop1K.   Copenhague

CoPer. Contr. Perkins Obs. CopRe. Copenhagen University Observatory,

Reprints

CoRad. Communications from the Radcliffe Ob-

servatory, Pretoria, South Africa

CoROE. Communications from the Royal Obser-

vatory, Edinburgh

CoRut. Contributions from the Rutherford Observatory of Columbia University, New York CoSka. Contr. Astron. Obs. Skalnate Pleso CoStA. Communications from the University

Observatory, ST Andrews, Scotland

CoSte. Contributions - University of Arizona,

Steward Observatory, Tucson, Arizona

CoTok. Contributions from the Department of

Astronomy - University of Tokyo

CoTol. Cerro Tololo Inter-amer. Obs. CoTor. Contrib. Oss. Astron. Torino CoUCL. Contributions - Universite Catholique de

Louvain

CoVVO. Contr. Van Vleck Obs. CoWat. Contr. Univ. Waterloo Obs. CoWro. Contrib. Wroclaw Astron. Obs. CR2... C.R. Acad. Sci. Ser. II CR.... C.R. Acad. Sci. CRA... C.R. Acad. Sci. Ser. A CraRe. Cracow Observatory - Reprints CRB... C.R. Acad. Sci. Ser. B CRJS3. C.R. sur les Journees de Strasbourg,

3eme Reunion

CSRV.. Catalog of stellar radial velocities

C&T...   Ciel et Terre

CVSS1. USSR Acad. Sciences,1973 (XXFX) CZC... R. Obs., Cape of Good Hope, London DDORe. David Dunlap Observatory, Richmond

Hill - Reprints

DGS75. La dynamique des galaxies spirales. Col-

loque international C.N.R.S. 241,pp

DOCi.. Documentation des Observateurs -

Circulaire

DOIAP. Doc. des Observ. Inst. d’Astr. de Paris DoSSR. Dokl. AN ArmSSR DudOR. Dudley Obs. Rep. DudRe. Dudley Observatory, Albany, New York

- Reprints

DunOP. Dunsink Observatory, Dublin, Ireland -

Publications

DunRe. Dunsink Obs. Rep. EBCi.. Eclipsing Binaries Circulars - Polska

Akademia Nauk, Krakow

EJPh.. European J. Phys. EL.... Europhysics Lett.

EM&P..   Earth

Endvr. Endeavour New Ser. ESO82. European Southern Obs. ESOAR. ESO Annu. Rep. ESOB.. European Southern Observatory -

Bulletin

ESOSP. European Southern Observatory - Scien-

tific Preprints

ESOSR. ESO Sci. Rep. ExA... Experimental Astron. EXOSA. EXOSAT Express FCPh.. Fundamentals of Cosmic Physics FoPh.. Found. Phys.

                                                GC....   GC

GCCat. Carnegie Inst. Washington D.C.

Publ.,1933 (XXAK)

GCCat. Carnegie Inst. Washington D.C.

Publ.,1937 (XXAL)

GCCCS. Publ. Warner & Swasey Obs. 3, No. 2 GCRV.. Carnegie Inst. Washington D.C. Publ.

601,1953 (XXBH)

GCTP2. Yale Univ. Obs., 1963,1963 (XXDX) GCTP.. Yale Univ. Obs. 1952,9,1952 (XXBF) GCVS1. Moscow, Acad. of Sciences USSR Shternberg,1951 (XXAN) GCVS2. GENERAL CATALOGUE OF VARI-ABLE STARS,1958 (XXCE) GCVS3. General Catalogue of Variable Stars, 3rd

ed.

Gemin. GEMINI Newsletter Royal Greenwich

Obs.

Gen1K. Observatoire de Geneve GeoOM. Georgetown Obs. Monogram GeoRe. Georgetown Observatory - Reprints

GEOSC.   GEOS Circ.

GEOSN. GEOS Note Circulaire - Group: Etude et Observation Stellaire, Paris GICi.. Gruppo Italiano RV Tauri - Circolare

(Massa)

GLORe. Goethe Link Observatory of Indiana University, Bloomington, Indiana - Reprints

GriO..   Griffith Obs.

Groom. Proceeding of the Royal Soc., Edinburgh GUL... Geschichte Lichtwechsel Verand. Sterne,

Potsdam

HalAR. Hale Observatories, Pasadena, Califor-

nia - Annual Report of the Director

HalRe. Hale Observatories, Pasadena, Califor-

nia - Reprints

HamS.. Hamburger Sternwarte - Sonderdrucke HarAC. Harvard College Obs. HarCi. Harvard Circ. HarRe. Harvard College Observatory - Reprints HarZi. Harthaer Beobachtungs-Zirkular HawRe. Reprints - University of Hawaii, Institute

for Astronomy

HD.... Ann. Astron. Obs. Harvard Coll. vol

91-99

HelOB. Helwan Institute of Astronomy and Geophysics (Helwan Observatory) - Bulletins

HelR.. Obs. Astrophys. Lab. Univ. Helsinki

Report

HelRe. Astronomical Observatory, Helsinki,

Finland - Reprints

HemD.. Hemel en Dampkring HvaOB. Hvar Obs. Bull.

IAL1K. Institut d’Astronomie de Lausanne

IAUC..   IAU Circ.

IAUT.. IAU Transactions IBSH.. Inf. Bull. South. Hemisph. IBVS.. IAU Inform. Bull. Var. Stars

Icar..   Icarus

IDS... Publ. Lick Obs. 21,part 1,1963 (XXDU) IIApN. Indian Inst. Astrophys. Newsletter IRA... Infrared Astronomy IrAJ.. Irish Astron. J. IRAS.. Joint IRAS Science,Working Group,1985

(XXIF)

IRC... Calif. Inst. Technology, NASA,1969

(XXFE)

ISKZ.. Issl. Solntsa Krasnykh Zvezd. IUEEN. IUE ESA Newsletter IUENN. IUE NASA Newsletter IzAne. Izvestiya na Fizicheskiya Institut s Aneb IzAsh. Izvestiya Akademii Nauk Turkmenskoj

SSR, Ashkhabad

IzKaz. Izvestiya Astronomicheskoj Engel’gardt

obskoj Observatorii, Kazan

IzKry. Izv. Krym. Astrofiz. Obs. IzmP.. Ege University, Izmir - Publications of

the Department of Astronomy

IzPul. Izv. Glav. Astron. Obs. JAF... J. Astronomes Francais JApA.. J. Astrophys. Astron. JASAC. Japan Astronomical Study Association -

Circulars

JASEg. J. Astron. Soc. Egypt JASV.. Journal of the Astronomical Society of

Victoria, Melbourne

JAVSO. J. Am. Assoc. Variable star obs. JBAA.. J. Br. Astron. Soc. JBIS.. J. Brit. Interpl. Soc. JBRe.. Astron. Contrib. Univ. Manchester, Jo-

drell Bank Repr

JHA... J. History of Astronomy

JO....   J. Obs.
JOp...   J. Opt.

JRASC. J. R. Astron. Soc. Can.

JRASN.   South. Stars

JWasA. Journal of the Washington Academy of

Sciences

KazOB. Astronomicheskoj Observatorii im V.P.

Engel’gardta, Kazan - Byulleten’

KFNT.. Kinematika Fiz. Nebesn. Tel. KodOB. Kodaikanal Obs. Bull. KodRe. Kodaikanal Obs. Repr.

Kozmo.   Kozmos

KPPH.. Astron. Inst. Basel KSVH.. Kungl. Svenska Vetensk. Hand. KVeBB. Kleine Veroeffentlichungen der Universi-

taetssternwarte zu Berlin- Babelsberg

Lalan. British Ass. Adv. Sci., London LASP.. Leaflet Astron. Soc. Pacific LAstr. L’astronomie LBcat. The Observatory, Univ. Minnesota, Minneapolis,1953 (XXCB) LCVS.. Lightcurve N.W.A.V.S.O.

(North Western Association of Variable Star Observers)

                                                Lei1K.   Leiden

LFT... Lund  Press,  Minneapolis,

Minnesota,1955 (XXCC)

LicOB. Lick Obs. Bull.

Lie57. 8eme Colloque Intern. d’Astrophys. a

Liege

LLS69. Low Luminosity Stars - Kumar Gordon.,

Breach LowOB. Lowell Obs. Bull. LP.... Univ. Minnesota, Minneapolis,$ fasc. 1-57,$ 1963-81,1963 (XXDV) LPlaS. Obs. Astron. La Plata - Sep. Astron. LS.... Hamburger Sternw., Warner & Swasey

Obs.

LTT... Lund Press, Minneapolis, Minnesota,1957 (XXCD) LTTN.. Lund Press, Minneapolis,

Minnesota,1961 (XXCP)

LTTS1. The Observatory, Univ. Minnesota, Minneapolis,1962 (XXDB)

Lub75. Proc. Southwest Region Conf.

LuyWD. Univ. Minnesota, Minneapolis,1970

(XXFK)

LyoRe. Observatoire de Lyon -Reprints

MeAar. Meddelser fra Ole Romer Observatoriet,

Aarhus

MeGen. Mededelingen - Universiteit te Gent,

Sterrenkundig Instituut

                                                MeLeu.   MKL

MeLu1. Lund Medd. Astron. Obs. Ser. I MeLu2. Lund Medd. Astron. Obs. Ser. II

Mercu.   Mercury

MerRe. Instituto Venezolano de Astronomia,

Merida, Venezuela

Meteo.   De Meteoor

Meu1K. Observatoire de Meudon MeUpp. Uppsala Astr. Obs. Medd. MiARB. Astron. Rechen-Inst. Heidelb., Mitt.,

Ser. B

MiARI. Mitt. Astron. Rechen-Inst. Heidelberg MiBon. Mitteilungen der Astronomischen Institute der Universitaet Bonn

MiBud. Mitt. Sternw. Ungar. Akad.

MicC.. Michigan catalogue

MicRe. Reprints from the Observatory of the University of Michigan, Ann Arbor

MiGra. Mitt. Universitatssternw. Graz

MiHam. Mitt. Hamburger Sternw.

MiHar. Mitteilungen der Bruno-H.-Buergel-Sternwarte, Hartha, DDR MiHei. Mitteilungen der Landessternwarte

Heidelberg-Koenigstuhl

MiInn. Mitteilungen der Sternwarte Innsbruck MiJen. Mitt. Univ. Sternw. Jena MiMue. Mitt. Sternw. Univ. Munster MiMun. Mitteilungen der Sternwarte Munchen MiPot. Mitteilungen des Astrophysikalischen

Observatoriums, Potsdam

MiSon. Mitt. der Sternwarte Sonneberg MitAG. Mitteil. Astron. Gesellschaft MiTau. Mitt. Karl-Schwarzschild Obs.

Tautenburg

MiTue. Mitteilungen des Astronomischen Insti-

tuts der Universitaet Tuebingen

MitVS. Mitt. Verand. Sterne MiWie. Mitt. Univ. Sternw. Wien MKAtl. Astrophys. monographs, Univ. Chicago

Press

MLS69. Mass Loss from Stars, ed. M. Hack MmARB. Mem. Acad. Roy. Belgique MmASI. Memoirs of the Astronomical Society of

India

MmBAA. Memoirs of the British Astronomical

Association

MmKyo. Mem. Coll. Sci. Univ. Kyoto MmMtS. Mem. of the Mount Stromlo Observ. MmRAS. Mem. R. Astron. Soc. MmSAI. Mem. Soc. Astron. Ital. MNRAS. Mon. Not. R. Astron. Soc. MNSSA. Mon. Notes Astron. Soc. South Afr. Msngr. The Messenger MSRSL. Mem. Soc. R. Sci. Liege MtSOM. Mount Stromlo Obs. Mimeo. MtSRe. Mount Stromlo Observatory, Canberra -

Reprints

MtWRe. Mount Wilson and Palomar Observatory

- Reprints

N30... N30 catalogue NaiRe. Reprints - Utta Pradesh State Observa-

tory, Naini Tal, Insia

NapCo. Osservatorio Astronomico di Capodimonte-Napoli - Contributi Astronomici

NASAC. National Aeronautics and Space Admin-

istration, Conference Publication

NASAR. NASA Ref. Publ. NASAS. NASA Spec. Publ. NATi.. Nordisk Astronomisk Tidsskrift

Natur.   Nature

NAZ... Nachrichtenblatt der Astronomischen

Zentralstelle, Heidelberg

NBSSP. Nat. Bur. Stand. Spec. Publ. NewSc. New Scientist Nic1K. Observatoire de Nice NInfo. Naucnye Informacii NizRe. Nizamiah Observatory, Hyderabad -

Reprints

NLTT.. University Minnesota- USA -,1986 (XXIJ)

NoDef.   Not Defined

NPhS.. Nature, Physical Science

NRAOR. National Radio Astronomy Observatory

- Reprints

NVS... Nachrichtenblatt der Vereinigung der

Sternfreunde e.V.

OBN1K. Observatoire de Besancon - CNES Obs... The Observatory OklRe. Reprints - University of Oklahoma

Observatory

ONRAS. Occasional Notes of the Royal Astro-

nomical Society

OOVRO. Observations Owens Valley Radio Obs.

                                                Orion.   Orion

ORROE. Occasional Reports R. Obs.

OslR.. Inst. Theor. Astrophys., Blindern-Oslo,

Rep.

OVS... Observation of Variable Stars

PA....   POP.ASTR.

PAAS.. Publ. Amer. Astron. Soc.

PAB... Publications of The Astrophysics Branch PadCR. Osservatorio Astronomico di Padova - Comunicazioni e Rassegne PAICz. Publ. Astron. Inst. Czech. Acad. Sci. PAllO. Publ. Allegheny PASAu. Proc. Astron. Soc. Aust. PASJ.. Publ. Astron. Soc. Jap. PASK.. Publ. Ac. Sc. Kasakstan, Sect.

Astrobotanics PASP.. Publ. Astron. Soc. Pac. PASPS. Publ. Astron. Soc. Pacific Suppl. PAth.. Publ. Labo. Astron., Univ. Athens PATi.. Popular Astronomisk Tidsskrift PAUTx. University of Texas - Publications in

Astronomy PAZh.. Pis’ma Astron. Zu. PBeiO. Publ. Beijing Astron. Obs. PBosO. Publ. Bosscha Obs. PBrn.. Astron. Inst. Univ. Brno, Publ. PCat.. Publ. Catania PCinO. Publ. Cincinnati Obs. PCopO. Publikationer og mindre Meddeler fra Kobenhavns Observatorium PDAO.. Publ. Dominion Astrophys. Obs.

Victoria

PDAUC. Publ. Dept. Astron. Univ. Chile PDDO.. Publ. David Dunlap Obs. PDO... Publications of the Dominion Observatory, Ottawa, Ontario, Canada

PGenA. Publ. Obs. Geneve Ser. A PGenB. Publ. Obs. Geneve Ser. B PGro.. Publ. Kapteyn Astron. Lab. Groningen Phoen. Phoenix, Mitteilungsblatt fuer Veraenderlichenbeobachter

PhR... Physics Reports PhRvA. Physical Review A PhRvB. Physical Review B PhRvC. Physical Review C PhRvD. Physical Review D PhRvL. Physical Review Letters PhS... Physica Scripta PhT... Physics Today PIAG.. Publications de l’Institut d’Astronomie

et de Geophysique, Louvain

PIstO. Publ. Istanbul Univ. Obs. PJA... Proceeding of the Japan Academy PKCat. Publ. House Czech. Acad. Sci.,1967

(XXEP)

PLicO. Publ. Lick Obs. PLPla. Obs. Astron. La Plata Ser. Astron. PN70.. Planetary Nebulae Reidel D. Publ.

Comp. Dordrecht

PNAOJ. Publ. Nat. Astron. Obs. Japan PNAS.. Proc. Nat. Acad. Sci. PoA... Postepy Astronomii, Krakow POAN.. Publ. Obs. Astron. Nacional, Cerro

Calan

POBeo. Publications de l’Observatoire

Astronomique de Beograd

POBol. Publ. dell’Osserv. Astron. Bologna POBor. Publications de l’Observatoire de

l’Universite de Bordeaux

POClu. Publ. Obs. Astron. Univ. Cluj POHel. Publications of the Astronomical Obser-

vatory, Helsinki, Finnland

POHP.. Publ. Obs. Haute-Provence POIst. Publ. Obs. Astron. Univ. Istambul POLyo. Publ. Obs. de Lyon POMic. Publ. Michigan Obs. POMil. Pubblicazioni dell’Osservatorio Astro-

nomico di Milano-Merate (Nuova Serie

POMin. Publ. Astron. Obs. Univ. Minnesota POPad. Publ. Oss. Padova POPal. Pubblicazioni dell’Osservatorio Astro-

nomico di Palermo

POPar. Publication de l’Observatoire de Paris -

Notes et Informations

POPot.   Post. Publ.

POStr. Publ. Obs. Strasbourg POTor. Pubbl. Varie Fuori Ser. Oss. Astron.

Torino

POTou. Publications de l’Observatoire de

Toulouse

POVRO. Publ. Owens Valley Obs. POxf.. Publications - University of Oxford, De-

partment of Astrophysics

PPMtO. Publ. Purple Mountain Obs.

Pram..   Pramana

PraP.. Astronomical Institute of the Charles

University - Publications

Priv.. Private Communication PROE.. Publ. R. Obs. Edinburgh PSAIL. Pubbl. Stn. Astron. Int. Latitudine,

Carloforte-Cagliari PSCDS. Publication speciale du C.D.S. PSci.. Pour la Science P&SS.. Plan. Space Sci. PTarO. Publ. Tartuskoj Astrofiz. Obs. PTRSL. Philos. Tr. R. Soc. London PUSNO. Publ. U.S. Naval Obs. Second Serie PW&SO. Publ. Warner & Swasey Obs. PYerO. Pub. Yerkes Obs. PZ.... Perem. Zvezdy PZP... Perem. Zvezdy Prilozh. QJRAS. Quart. J. R. Astron. Soc. RA.... Ricerche Astron. RadRe. Radcliffe Observatory, Pretoria, South

Africa - Reprints

Rech.. La Recherche ReToh. Sci. Reports Tohoku Univ. Ser. I RGOB.. R. Obs. Bull. RioB.. Observatorio Nacional Publicacoes do

Servico Astronomico, Rio de Janeiro

Rise..   Rise hvezd

RMxAA. Rev. Mex. Astron. Astrofis. RNAO.. Results of National Argentine Obs. RoAJ.. Romanian Astronomical Journal

ROAn..   R. Obs. Ann.

ROCi.. Republic Obs. Circ. Roczn. Rocznik Astronomiczny Observatorjum

Krakowskiego, Krakow

ROLun. Reports Lund Obs. RomCo. Osservatorio Astronomico di Roma su

Monte Mario - Contributi Scientifici

RosOB. Observatorio Astronomico Municipal de

Rosario, Argentia - Boletin

RPPh.. Rep. Prog. Phys.

RvA...   Rev. Astron.

RvMad. Revista Real Acad. Ciencias Exact.,

Fis., Nat. Madrid

RvPD.. Revue du Palais de la Decouverte

RZh...   R.J.UDSSR

SAAOC. South Afr. Astron. Obs. Circ. SAnAp. Ann. Astrophys. Suppl. SanS.. Univ. de Chile, Depart. de

Astron.Separata

SAO... SMITHSONIAN INSTITUTION,

WASHINGTON DC,1966

SAO... SMITHSONIAN INSTITUTION,

WASHINGTON DC,1966 (XXEL)

SAOSR. SAO Special Report SASn.. Studia Astronomica Sinica SCA... Studii si Cercetari de Astronomie,

Bucuresti

Sci...   Science

SciAm. Scientific American

SciN..   Science News

SenAR. Sendai Astron. Rap. SFSN.. Scripta Fac. Sci. Nat. Ujep Brunensis

Physica

SoByu. Soobshch. Byurakan Obs.

SoKie. Sonderdrucke der Sternwarte Kiel SoMue. Sonderdrucke - Universitaet Muenster,

Astronomisches Institut

SoSAO. Soobshch. Spets. Astrof. Obs. SoSht. Soobshch. Gos. Astron. Inst. P.PK.

Shternberga

SPh... Solar physics SPhD.. Soviet Phys. Doklady SprRe. Sproul Observatory - Swarthmore, Penn-

sylvania - Reprints

SRToh. Sci. Rep. Tohoku Univ., Eighth Ser. SSRv.. Space Sci. Reviews SSTor. Std. Soc. Sc. Torunensis

S&T...   Sky Telesc.

StARe. Reprints - University Observatory, St.

Andrews, Scotland

Sterb.   Sternenbote

SteRe. Reprints - University of Arizona, Stew-

ard Observatory, Tucson, Arizona -

Stern.   Die Sterne

StoAn. Stockholm Obs. Ann. StoMe. Stockholms Observatoriums -

Meddelande

StoOR. Stockholms Obs., Rep. SUVSR. Scandinavian Union of Amateur As-

tronomers, Variable Star Section - Reports

SvA...   Sov. Astron.

SvAL.. Soviet Astron. Lett. SvPhA. Soviet Physics-Astronomy S&W... Sterne und Weltraum SydOP. Sydney Obs. Pap.

Tac1K.   Tache 1000

TamCo. Astronomical Contributions from the

University of South Florida at Tampa

TApA.. Top. Astrophys. Astron. Space Sci. TAPS.. American Philosophical Society TarOT. Tartu Astrofuus. Obs. Teated TD1... TD1 satellite TerCo. Osservatorio Astronomico di Collurania,

Teramo - Contributi

TerMm. Osservatorio Astronomico di Collurania,

Teramo - Memorie ed Osservazioni

TerNC. Osservatorio Astronomico di Collurania,

Teramo - Note e Comunicazioni

Thesi.   Thesis

TITas. Tsirk. Astron. Inst. Tashkent TMSS.. Two micron sky survey TokAB. Tokyo Astron. Bull. TokRe. Tokyo Astronomical Observatory -

Reprints

TOYal. Trans. Astron. Obs. Yale Univ. TrAlm. Tr. Astrofiz. Inst. Alma-Ata TrDus. Trudy Instituta Astrofiziki, Dushanbe -

Akademiya Nauk Tadzhikskoj SSR

TreP.. Osservatorio Privato Specola ”Ariel” Treviso Italia - Pubblicazione TriP.. Observ. Astr. Trieste TrKaz. Tr. Kazan. Gorod. Astron. Obs. TrLen. Trudy Astron. Obs. Leningrad TrPul. Tr. Glav. Astron. Obs. Pulkovo, ser.

2,pp

TrRig. Trudy Astrofiz. Lab. Riga

TrSht. Trudy Gosud. Astron. Sternberga TrSta. Trudy Instituta Astrofiziki, Stalinabad - Akademiya Nauk Tadzhikskoj SSR TsPul. Tsirkulyary Glavnoj Astronomicheskoj

Observatorii i Pulkove TsShe. Tsirk. Shemakh. Astrofiz. Obs. TSSLW. Travaux de la Societe des Sciences et des

Letters de Wroclaw

UCLAP. Astronomical Papers - University of Cal-

ifornia, Los Angeles

UGC... Nova Acta Regiae Soc. Sci. Upsaliensis,

Ser. V:A

ULIA4. Univ. de Liege, Inst. Astrophys.

ULIA8. Univ. de Liege, Inst. Astrophys. Coll.

in 8 No

UMt1K. Universite de Montpellier Unpub. Unpubl.,1968 (XXEU) UppAn. Uppsala Astron. Obs. Ann. UppOR. Upps. Astron. Obs. Rep.

                                                Urani.   Urania

USNOC. Circulars - U.S. Naval Observatory,

Washington D.C.

USNOR. Reprint - U.S. Naval Observatory, Wash-

ington D.C.

UtrOv. Utr. Sterrekundige Overdrukken VA.... Vistas in Astronomy VAG... Vierteljahresschrift der Astronomischen

Gesellschaft

VatAR. Specola Astronomica Vaticana - Annual

Reports

VatCo. Specola Astronomica Vaticana -

Comunicazione

VatMA. Specola Astronomica Vaticana - Miscel-

lanea Astronomica

VatOP. Vatican Obs. Publ.

VatRA. Specola Astronomica Vaticana -

Richerche Astronomiche VeARI. Veroeff. Astron. Rechen-Inst. VeBB.. Veroeffentlichungen der Universi-

taetssternwarte zu Berlin-Babelsberg

VeBoc. Veroeffentlichungen des Astronomischen

Instituts der Ruhr-Universitaet, Bochum

VeBon. Veroeff. Astron. Inst. Bonn

VeGoe. Veroeffentlichungen der Universitaets-Sternwarte zu Goettingen VeHei. Veroeff. des Badischen Sternw.

Heidelberg VeMun. Veroeff. Munchen VeSon. Veroeff. Sternwarte Sonneberg VeUSB. Veroeff. Univ. Sternw. Bonn VeWFS. Veroeffentlichung der Wilhelm Foerster

Sterwarte VilCo. Villanova Univ. Obs. Contr. VilOB. Vilnius Astron. Obs. Biul. VKha.. Vestnik Khar’kovskogo Universiteta VKie.. Vestn. Kiev. Univ. Astron. Vsin.. Contr. Oss. Astrof. Univ. Padova in

Asiago 239,250,294

VSSCi. Royal Astronomical Society of New Zealand, Variable Star Section - Circulars

VSSRe. Royal Astronomical Society of New

Zealand, Variable Star Section - Reprints

VVORe. Reprints - Van Vleck Observatory, Mid-

dletown, Connecticut

WroRe. Wroclaw Astron. Obs. Repr. YamCi. Yamamoto Circ. ZA.... Zeitschr. Astrophys. ZD.... Zvaigznota Debess

Zenit.   Zenit